OpenStudy (anonymous):

A cargo plane flew to the maintenance facility and back. It took one hour less time to get there than it did to get back. The average speed on the trip there was 220mph. The average speed on the way back was 200mph. How many hours did the trip there take?

OpenStudy (anonymous):

Set up the problem: x = time it took to get there in hours y = time it took to get back in hours x +1 = y 220x = 200y solve now 220x = 200(x+1) 220x=200x+200 20x = 200 x= 10